OpenStudy (anonymous):

The area of a rectangle is 8x^2+14x+3. The length is 2x+3. what is the width? PLEASE SHOW WORK! : )

OpenStudy (anonymous):

a=l*b b=a/l

OpenStudy (anonymous):

Can you do it?

OpenStudy (anonymous):

\[8x^2+14x+3\over 2x+3\]

OpenStudy (anonymous):

\[8x^2+7\]

OpenStudy (anonymous):

Thanks So Much!!!!!!!!!!1111

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

Would you be able to help me with my next problem?

OpenStudy (anonymous):

ya say @LanaVon

OpenStudy (anonymous):

Ok. Gimme some time to write it

OpenStudy (anonymous):

The surface area of a cylinder can be found using the expression 2piR^2+2piRH, where r represents the radius of the cylinder and h height. Factor the expression

OpenStudy (anonymous):

@Rohangrr